OpenStudy (anonymous):

Find the square root of 17,424 without using calculator or tables

OpenStudy (saifoo.khan):

Prime Factors!

OpenStudy (bahrom7893):

hmm try dividing this by 121

OpenStudy (amistre64):

with squareroots; divvy it up into pairs of 2 from the decimal like this: 1 74 24 now we can slice and dice it

OpenStudy (bahrom7893):

Wait or would that be considered using a table?

OpenStudy (anonymous):

a64 yes that is what I need

OpenStudy (amistre64):

what is the closest square that equals "1" ?

OpenStudy (amistre64):

by hand, we address x^2 + 2xy + y^2 format

OpenStudy (anonymous):

use long division

OpenStudy (amistre64):

we need to start with an x^2 that is close to, but not greater then, our target grouping. in this case, its a "1" :)

OpenStudy (bahrom7893):

I would use long division to divide 17,424 by 121 since 121 is 11^2, but i think that would count as memorizing the table

OpenStudy (amistre64):

the "by hand" steps weed out any guess work

OpenStudy (bahrom7893):

i c

OpenStudy (amistre64):

our first target group in this case is "1"; and we wish to find an "x^2" that is close to or equal to it. 1^2 = 1 ... thats great :) 1-1 = 0; now drop down the next grouping, in this case its 74 0 74 ; the zero is there just for show; lets focus on the 74 part.. this next step is the "tricky" part; we know our x^2, and we gotta find a (2x)y that matches our grouping. ... recall that x=1 in this case x y ----- ; this is our setup 2xy | 074 ? ----- 2(1)__ | 074 1 ? ----- ; twenty "something" times "the unknown" is equal 2__ | 074 to or less then "74" 20*0 = 0 21*1 = 21 22*2 = 44 23*3 = 69 24*4 = 96 .... to big, lets go with 3 1 3 ----- 23 | 074 -69 ----- 5 ; now attach our next grouping (24) to this and repeat 13 ? ; since x = 13 now, (2x)y is 26__ ----- 26__ | 524 260*0 = 0 261*1 = 261 262*2 = 524 .. that one works 13 2 ; since x = 13 now, (2x)y is 26__ ----- 262 | 524 524 ----- 0 the answer is; 132

OpenStudy (anonymous):

Thanks

OpenStudy (amistre64):

if any of that doesnt make sense, let me know so I can try to clarify it :)

OpenStudy (anonymous):

It does :D

OpenStudy (amistre64):

great :)

OpenStudy (amistre64):

the cubic are much worse ... and dont even get me started and 4roots lol